In a system a particle A of mass m and charge -2q is moving in the nearest orbit around a very heavy particle B having charge +q .Assuming Bohr's model of the atom to be applicable to this system the orbital angular velocity of the particle A is

Answer: (2πmq^4)/(€²h^3)
1
Expert's answer
2020-02-26T10:24:52-0500

"F=ma"


"\frac{2q^2}{4\pi \epsilon_0r^2}=m\frac{v^2}{r}\to m^2v^2=\frac{2mq^2}{4\pi \epsilon_0r}"


According to Bohr's model of the atom


"mvr=n\frac{h}{2\pi}". In our case "n=1"


"\frac{2mq^2}{4\pi \epsilon_0r}=\frac{h^2}{4\pi^2r^2}\to r=\frac{\epsilon_0h^2}{2\pi mq^2}"


"v=\frac{h}{2\pi mr}=\frac{2\pi hmq^2}{2\pi m\epsilon_0h^2}=\frac{q^2}{\epsilon_0h}"


So, we have


"v=\omega r\to \omega=\frac{v}{r}=\frac{q^2}{\epsilon_0h}\cdot \frac{2\pi mq^2}{\epsilon_0h^2}=\frac{2\pi mq^4}{\epsilon^2_0h^3}" Answer.